OpenStudy (anonymous):

Given that i^2 =-1 compute the following powers of i. Please show steps. i^3

OpenStudy (anonymous):

\[i = \sqrt{-1}\] \[i ^{3} = i ^{2}*i = -1*\sqrt{-1} = -\sqrt{-1}\]

OpenStudy (anonymous):

\[i^0=1,i^1=i, i^2=-1, i^3=-i, i^4=1,\] lather, rinse, repeat